Prev Next

SW&IT Meeting Minutes 2021-03-22

Attendees: David, Igor, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210308
LR20210308 - Bec, Matthieu 80.0
LR20210308 - Edmeades, David 82.0
LR20210308 - Sola, Igor 0.0
LR20210308 - Sizemore, Michael 13.0
LR20210308 - Grenz, Paul 20.0
LR20210308 - Hooper, Stephen 80.0
LR20210308 - Zhang, Yang 70.0
LR20210308 - Zhang, Xianyu 64.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management
    • ATT PM priorities, MCS, candidates review/interview
    • OT for SHARKs
    • Review storage document
  • This week
    • ATT PM FTE & priorities, due next week
    • candidates review

YZ

  • Previous:
    • LBC stuff (IT8329 RB_Science, and IT8348, Blue Profibus).
    • MODS IT8345 (acqMODS bugs), now closed.
    • TCS support call (PCSGUI trouble, Stephen helped with tcs1/2 reboot/glusterfs mount, also some zoom discussion with Stephen).
    • TMS software next phase discussion (mainly with John Hill).
    • Some Web-dms dev.
    • Some TMS utility dev.
    • OSCO abandoned on Saturday and Sunday night.
  • Next:
    • Release a new RB_Science.
    • Upgrade TMS utility, plus some wiki editing work.
    • Web-dms dev.
    • TMS mountain mongodb setup? and experiment.
    • OSCO meeting (and hopefully new data).
    • Maybe restart MODS 64-bit work.
    • Safety training overdue!!

DE

  • Synology prep, office trip
    • Did not go well
    • Waiting on Syno support for sanity check and instructions specific to our situation. Likely another office trip this week
  • Storage work
  • Cloud storage meetings
  • Veeam quote
Next:
  • Bother Synology
  • More meetings
  • Fix Synology
    • Support is offering to remote in to fix the versioning, but will result in downtime. Thoughts on timing? I'm not sure if weekends are available, but can check if that's what we decide.
  • Safety training

XZ

  • Last week
    • ATT AdSec tasks
    • AO support
    • Document for AO SW configuration and release control management
    • Thursday and Friday is out of office
  • This week
    • ATT AdSec /SW
    • DSPCoil Current issue
    • Safety training
    • AO support

PSG

  • Last week
    • Looking at OAC code
    • Looking into LBTI Website access.
  • This week
    • Set up time to go into the lab and test UMAC

SH

  • VM stuff (little more progress)
  • DIMM with LeRoy
  • TCS incident, meeting
  • John Hill VPN and e-mail
  • obs issues with AD
  • Network switch UITS
  • IssueTrak stuff for Laura
  • UA Secure

next:

  • begin phase II move to LDAP
  • some more VM stuff
  • DIMM with LeRoy (maybe)
Topic revision: r8 - 22 Mar 2021, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback